OpenStudy (cutiekawdfi):

help please (: fan and medal find the number of roots for each equation. 1. 2 - x^4 + x^2 = 0 explain pls <3

OpenStudy (anonymous):

Step 1: Write the equation in standard form

OpenStudy (danjs):

most likely you can factor this , if it is a 'nice' book problem

OpenStudy (danjs):

it is order 4 from the x^4 term, it can have up to 4 solutions to that , zeroes, real or complex

OpenStudy (danjs):

Can you factor x^2 - x - 2 ?

OpenStudy (cutiekawdfi):

(x-2)(x+1)

OpenStudy (cutiekawdfi):

@DanJS

OpenStudy (cutiekawdfi):

@DanJS @DanJS @DanJS

OpenStudy (anonymous):

This equation can be re-written as: x^4 - x^2 +2 = 0 Plug y = x^2

OpenStudy (danjs):

good factoring, think of x^4 as [x^2]^2 and x^2 as [x]^2, lik telltoamit says, substitute u = x^2 u^2 - u + 2 = 0 you factored that to (u - 2)(u + 1) = 0 now replace back the x^2 from the substitution u=x^2, and see what that does

OpenStudy (danjs):

(x^2 - 2) * (x^2 + 1) = 0 you can solve both those quantities and get 2 real and 2 complex roots

OpenStudy (cutiekawdfi):

is tht the answer

OpenStudy (danjs):

no, remember solving these, if one quantity is zero, then the equation will be true, so you can solve x^2 - 2 = 0 or x^2 + 1 = 0 each of those will give you 2 answers

OpenStudy (danjs):

x^2 = 2 \[x = \pm \sqrt{2}\] x^2 = -1 \[x = \pm \sqrt{-1} = \pm i\] look good?

OpenStudy (cutiekawdfi):

yea

OpenStudy (danjs):

most the time, if you have a higher order polynomial like that and they want you to solve it, most likely it will factor someway, if it is a book prob

OpenStudy (cutiekawdfi):

is that the answer

OpenStudy (danjs):

it asks for the number of roots... roots means when the polynomial y=p(x) = 0 , is zero, when y=0, or the x-intercepts.... here you see it has 4 answers, in general, a polynomial of degree 'n' will have at most 'n' zeroes.

OpenStudy (danjs):

sorry, 'zeroes' and 'roots' mean the same

OpenStudy (cutiekawdfi):

im soncused

OpenStudy (danjs):

here is a thing you can look over http://cs.franklin.edu/~sieberth/MATH160/bookFiles/Chapter3/333371_0302_263-275.pdf